if 4.695 g of pure element contains 4.449x10^22 atoms, what is the element?
1
Expert's answer
2020-09-14T08:25:53-0400

The molecular mass of the element is unknown hence we let it be "(xg\/mol)"

The number of moles of the element "x" is also unknown hence "(ymol)"

Converting the elements from Mass"\to" Moles"\to" Atoms gave us "4.449\times 10^{22}" atoms of "x"

In finding the molecular weight of the unknown element, we convert from

Atoms"\to" Moles"\to" Mass

Atoms of element "x" "=4.449\times10^{22}"

Moles of the element "x" (using the Avogadro Constant)

"6.022\times 10^{23}x\over 1mol" "\times ymol x" "=4.449\times 10^22 atoms x"

"y=0.0739mol x"

The number of moles of this element is thus "0.0739 mol es"

Converting Moles "\to" Mass

"4.695gx\times" "1mol x\over xg" "=0.0739mol x"

"x=63.53g"

The molecular weight of the element "=63.53g"

From the periodic table the element is Copper(Cu)